OpenStudy (anonymous):

bruce discovered that the wishing well contained 9 dollars in nickels and dimes. and that there were 30 more dimes than nickles. how many coins of each type were there?

OpenStudy (anonymous):

i corrected the bus problem - sorry about that

OpenStudy (anonymous):

before i do this - im from uk - i dont know what a dime or a nickel is!

OpenStudy (alfie):

0.5n + 0.10n = 9 0.5n + 0.10(n+30) = 9 Solve for n, you have n= 10 So you have 10 nickels and 40 dimes. You can find if it is correct with: 0.5(10) + 0.10(40) = ? if it is equals to 9 then it is correct.